If there are 8.40 moles of O, how many moles of each of the compounds are present?

H2SO4

mol

C2H4O2

mol

NaOH
Answer #1

1)

1 mol of H2SO4 has 4 moles of O

So,

Moles of H2SO4 = moles of O / 4

= 8.40 / 4

= 2.10 mol

Answer: 2.10 mol

2)

1 mol of C2H4O2 has 2 moles of O

So,

Moles of C2H4O2 = moles of O / 2

= 8.40 / 2

= 4.20 mol

Answer: 4.20 mol

3)

1 mol of NaOH has 1 mol of O

So,

Moles of NaOH = moles of O / 1

= 8.40 / 1

= 8.40 mol

Answer: 8.40 mol